A CD starts from rest and accelerates to an angular frequency of 3.5 rev/s. Calculate

the disc s average period T and centripetal velocity at the edge of the disc when

radius is 2. 0 x 10-2m. 


1
Expert's answer
2022-04-27T18:24:00-0400

"T=\frac 1\nu=0.29~s,"

"v=\omega r=2\pi \nu r=0.44~\frac ms."